### v2

@@ Appendix C Risk-seeking Policy Gradient
When using the empirical Lyapunov risk in Equation (4) as reward function, the risk-seeking policy gradient is not guaranteed to be unbiased.
@@
For a given dataset $\mathcal{X}\subseteq\mathcal{D}$, the empirical Lyapunov risk is a random variable whose value is determined by the random expression sampled from the symbolic transformer model. Let $Z$ denotes the negated empirical Lyapunov risk, with its probability density $p(z|\phi)$ depending on transformer parameters $\phi$.
@@
To effective penalize the invalid sampled expressions which miss some state variables or is symbolically invalid, their empirical Lyapunov risks are set to be $\infty$, or a symbolically valid candidate which incorporates all state variables might be less preferable than a symbolically invalid or variable-incomplete expression during the training. Consequently, $s_{\alpha}$ has a non-zero probability density of being $\infty$- ∞. If $s_{\alpha}=-\infty$, then the gradient of the expectation in Equation (8) is:
@@
Since the integration lower-bound is $\infty$- ∞, the Leibniz rule does not apply, which means interchange the gradient and the integration changes the result. Therefore, the policy gradient in Equation (8) is no longer guaranteed to be unbiased. ∎
@@
Reward Calculation. The reward function in Equation (5) quantifies the degree of violation based on empirical data points. However, as the dimensionality of the input dynamics increases linearly, the training set $\mathcal{X}$ must grow exponentially to maintain precision in the empirical measurement of violation degree, which is impractical for high-dimensional cases. To ensure the reward signal remains a reliable indicator of expression quality in the risk-seeking policy gradient without requiring an excessively large dataset, we incorporate Projected Gradient Descent (PGD) (Madry et al., 2017) into the reward calculation process. Prior to evaluating the reward, PGD is employed on all sampled expressions to efficiently identify a set of “minimizers” in the state space. These minimizers are computed from a randomly sampled set of starting points through PGD for each sampled expression, and the process can be parallelized on GPU, enabling efficient computation. Though PGD may converge to local minima, it remains effective in identifying a few counter-examples that violate Lyapunov conditions across invalid candidates. These counter-examples are then added to the training set $\mathcal{X}$ for reward calculation of all sampled expressions in the current epoch and are removed immediately after the calculation. This process leverages PGD to capture violation data points in advance, ensuring the quality of the reward signal without the need for an excessively large dataset.
@@ Appendix D Genetic Programming
In our framework, we employ three evolutionary operations: mutation, crossover, and selection, within our Genetic Programming component (DEAP (Fortin et al., 2012)). A mutation operator introduces random variations to an expression, such as replacing a subtree of one expression with another randomly generated subtree. A crossover operator exchanges content between two expressions, e.g. by swapping a subtree of one expression with a subtree of another expression, enabling the combination of their features. A selection operator determines which expressions persist onto the next population. A common method is tournament selection (Koza, 1992), where a set of $l$ candidate expressions are randomly sampled from the population, and the expression with the highest fitness value is selected. In each iteration of GP evolution, each expression has a probability of undergoing mutation and a probability of undergoing crossover; selection is performed until the new generation’s population has the same size as the current generation’s population. In empirical experiments, we set the probability of undergoing mutation and crossover to be $0.5$, and we adjust the size of the tournament and number of evolutions proportional to the dimension of the input system.
@@ E.1 Augmented Neural Lyapunov Control (ANLC)
The Augmented Neural Lyapunov Control (ANLC) (Grande et al., 2023) combines Artificial Neural Networks (ANNs) with Satisfiability Modulo Theories (SMT) solvers to synthesize stabilizing control laws to the input dynamics $f(x)$ with formal guarantee. The neural network is trained over a dataset of state-space samples to generate candidate control laws and Lyapunov functions, while the SMT solvers are tasked with certifying the Lyapunov conditions of the neural Lyapunov function over a continuous domain and returning a counterexample if function is invalid. To ease the computationally-inefficient verification process in SMT module, ANLC proposed a discrete falsifier, which discretized the state space for sample selection and evaluation, employed before the SMT call to avoid the frequent calling of time-consuming SMT falsifier. As the previous learning-based Lyapunov function construction approaches usually initialized the parameters of control policy with pre-computed gains from state-feedback controllers, e.g. Linear-Quadratic Regulators, which requires user time and control expertise to properly perform the initialization process, ANLC instead removes the need of control initialization by its proposed compositional control architecture containing both linear and nonlinear control laws so that the proposed method allows the synthesis of nonlinear (as well as linear) control laws with the sole requirement being the knowledge of the system dynamics. For empirical experiments, we tested the ANLC algorithm for all system dynamics in Appendix F and G. We tested on Van Der Pol Oscillator and 3D Trig dynamics to get the best hyperparameter setting. In bold we show the chosen parameters, selected to have the best success discovery rate on Van Der Pol Oscillator and 3D Trig Dynamics.
@@
FOSSIL 2.0 (Edwards et al., 2024) is a software tool for robust formal synthesis of certificates (e.g. Lyapunov and barrier functions) for dynamical systems modelled as ordinary differential and difference equations. FOSSIL 2.0 implements a counterexample-guided inductive synthesis (CEGIS) for the construction of certificates alongside a feedback control law. In the loop of CEGIS, the learner, based upon neural network templates, acts as a candidate to satisfy the conditions over a finite set $\mathcal{D}$ of samples, while the verifier (formal verification tools) works in a symbolic environment that either confirms or falsifies whether the candidate from learner satisfies the conditions over the whole dense domain $\mathcal{X}$. If the verifier falsifies the candidate, one or more counter-examples identified from the verifier are added to the samples set, and the network is retrained. This loop repeats until the verification proves that no counter-examples exist or until a timeout is reached. Similar to the ANLC, in the empirical experiment, we set the hyperparameters based on the Van Der Pol Oscillator and 3-D Trig dynamics and tested for all other dynamics in Appendix F and G.
@@ F.3 Three-variable-polynomial-systems with higher degree
Table 3: Three Dimensional Polynomial Example with Higher Degree.
@@
Our proposed method is trained over the state space $\mathcal{D}=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{6}\bm{\mid}|x_{i}|\leq 1,\forall i\in\{1,2,..,6\}\}$, and is able to find a valid Lyapunov function $V(x)=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}$. This dynamics is not globally asymptotically stable since if $x_{1},x_{2},\text{ or }x_{5}$ has a significant large value, the perturbations introduced by the small gains will shift the object by a significant amount away from the equilibrium point. By empirical checking, our found Lyapunov function certifies the asymptotical stability of this system over the region $\mathcal{D}^{\prime}=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{6}\bm{\mid}\sum_{i=1}^{6}x_{i}^{2}\leq 500\}$.
@@ G.3 N-bus Lossless Power System
We test our proposed framework on the N-bus power lossless system Cui et al. (2023a) to examine its ability on complex high-dimensional dynamics. Consider $\theta_{i}$, $\omega_{i}$ as the phase angle and the frequency of bus $i$, respectively, the dynamics for each bus are formulated as follows,
@@
where $\delta_{i}$ can be understood as the center-of-inertia coordinates of each bus. In our experiment, we test proposed frameowrk on 3-bus power system. For simplicity, we set $p_{i}=0$, $m_{i}=2$, $d_{i}=1$, $u_{i}(\omega_{i})=\omega_{i}$, and $B_{ij}=1\>\forall\>i\neq j,B_{ii}=0$. In this case, the equilibrium point for our system is at the origin, i.e. $\delta_{i}^{*}=\omega_{i}^{*}=0,\>i=1,2,3$. The state space for our experiment is defined as: $\mathcal{D}=\{(\delta_{1},\delta_{2},\delta_{3},\omega_{1},\omega_{2},\omega_{3})\in\mathbb{R}^{6}\bm{\mid}|\delta_{i}|\leq 0.75\text{ and }|\omega_{i}|\leq 1.2\text{ for }i=1,2,3\}$. Through our method, we retrieved a valid Lyapunov function $V(\delta_{1},\delta_{2},\delta_{3},\omega_{1},\omega_{2},\omega_{3})=(\sum_{i=1}^{3}\omega_{i}^{2})-0.5\left(\sum\limits_{i=1}^{3}\sum\limits_{\begin{subarray}{c}j=1,i\neq j\end{subarray}}^{3}\cos(\delta_{i}-\delta_{j})-1\right)$, which is consistent to the known Lyapunov function presented in Cui et al. (2023a). The lie derivative of identified Lyapunov function can be simplified as $L_{f}V=-2(\omega_{1}^{2}+\omega_{2}^{2}+\omega_{3}^{2})$. The analytical structure of this found Lyapunov function and invariance principle allow us to easily identify it as a valid Lyapunov function by hands.
@@ G.4 Indoor Micro Quadrotor
For the angular rotations subsystems of quadrator from Bouabdallah et al. (2004), it has 6 states to describe the angular motion of quadrotor. The states $x_{1},x_{3},\text{ and }x_{5}$ describes the roll, pitch, and yaw of the quadrator, and states $x_{2},x_{4},\text{ and }x_{6}$ represents their time derivatives. With perturbation terms $\Omega$ and control inputs $U_{1},U_{2},\text{ and }U_{3}$, the subsystem can be formulated as follows,
@@
and restricting $I_{x}=I_{y}$, the angular rotations subsystems is stabilized to the chosen equilibrium point $X_{d}=\{x_{1}^{d},0,x_{3}^{d},0,x_{5}^{d},0\}$. In empirical experiments, we set $X_{d}=\{0,0,0,0,0,0\}$, state space $\mathcal{D}=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{6}\bm{\mid}|x_{i}|\leq 3,\forall i\in\{1,2,..,6\}\}$, $I_{x}=I_{y}=2$, $I_{z}=5$, $k_{1}=5$, $k_{2}=20$, $k_{3}=4$, $l=1$, $J_{R}=1$, and $\Omega=\sin(x_{2})\cos(x_{4})$, our framework successfully found Lyapunov function $V(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=\sum_{i=1}^{6}x_{i}^{2}$. By examining the lie derivative $L_{f}V=-2.5x_{2}^{2}-10x_{4}^{2}-0.8x_{6}^{2}$ and using Invariance principle, we conclude that this subsystem is globally asymptotically stable. Under other parameter settings, our framework can also retrieve valid analytical Lyapunov functions for this dynamics.

### v3

@@ Appendix C Risk-seeking Policy Gradient
When using the empirical Lyapunov risk in Equation (4) as the reward function, the risk-seeking policy gradient is not guaranteed to be unbiased.
@@
For a given dataset $\mathcal{X}\subseteq\mathcal{D}$, the empirical Lyapunov risk is a random variable whose value is determined by the random expression sampled from the symbolic transformer model. Let $Z$ denote the negated empirical Lyapunov risk, with its probability density $p(z|\phi)$ depending on transformer parameters $\phi$.
@@
To effectively penalize the invalid sampled expressions that miss some state variables or are symbolically invalid, their empirical Lyapunov risks are set to be $\infty$, or a symbolically valid candidate that incorporates all state variables might be less preferable than a symbolically invalid or variable-incomplete expression during the training. Consequently, $s_{\alpha}$ has a non-zero probability density of being $\infty$- ∞. If $s_{\alpha}=-\infty$, then the gradient of the expectation in Equation (8) is:
@@
Since the integration lower-bound is $\infty$- ∞, the Leibniz rule does not apply, which means interchanging the gradient and the integration changes the result. Therefore, the policy gradient in Equation (8) is no longer guaranteed to be unbiased. ∎
@@
Reward Calculation. The reward function in Equation (5) quantifies the degree of violation based on empirical data points. However, as the dimensionality of the input dynamics increases linearly, the training set $\mathcal{X}$ must grow exponentially to maintain precision in the empirical measurement of violation degree, which is impractical for high-dimensional cases. To ensure the reward signal remains a reliable indicator of expression quality in the risk-seeking policy gradient without requiring an excessively large dataset, we incorporate Projected Gradient Descent (PGD) (Madry et al., 2017) into the reward calculation process. Prior to evaluating the reward, PGD is employed on all sampled expressions to efficiently identify a set of “minimizers” in the state space. These minimizers are computed from a randomly sampled set of starting points through PGD for each sampled expression, and the process can be parallelized on a GPU, enabling efficient computation. Though PGD may converge to local minima, it remains effective in identifying a few counterexamples that violate Lyapunov conditions across invalid candidates. These counter-examples are then added to the training set $\mathcal{X}$ for reward calculation of all sampled expressions in the current epoch and are removed immediately after the calculation. This process leverages PGD to capture violation data points in advance, ensuring the quality of the reward signal without the need for an excessively large dataset.
@@ Appendix D Genetic Programming
In our framework, we employ three evolutionary operations: mutation, crossover, and selection, within our Genetic Programming component (DEAP (Fortin et al., 2012)). A mutation operator introduces random variations to an expression, such as replacing a subtree of one expression with another randomly generated subtree. A crossover operator exchanges content between two expressions, e.g., by swapping a subtree of one expression with a subtree of another expression, enabling the combination of their features. A selection operator determines which expressions persist into the next population. A common method is tournament selection (Koza, 1992), where a set of $l$ candidate expressions is randomly sampled from the population, and the expression with the highest fitness value is selected. In each iteration of GP evolution, each expression has a probability of undergoing mutation and a probability of undergoing crossover; selection is performed until the new generation’s population has the same size as the current generation’s population. In empirical experiments, we set the probability of undergoing mutation and crossover to be $0.5$, and we adjust the size of the tournament and number of evolutions proportional to the dimension of the input system.
@@ E.1 Augmented Neural Lyapunov Control (ANLC)
The Augmented Neural Lyapunov Control (ANLC) (Grande et al., 2023) combines Artificial Neural Networks (ANNs) with Satisfiability Modulo Theories (SMT) solvers to synthesize stabilizing control laws for the input dynamics $f(x)$ with formal guarantees. The neural network is trained over a dataset of state-space samples to generate candidate control laws and Lyapunov functions, while the SMT solvers are tasked with certifying the Lyapunov conditions of the neural Lyapunov function over a continuous domain and returning a counterexample if the function is invalid. To ease the computationally inefficient verification process in the SMT module, ANLC proposed a discrete falsifier, which discretized the state space for sample selection and evaluation, employed before the SMT call to avoid the frequent calling of the time-consuming SMT falsifier. As the previous learning-based Lyapunov function construction approaches usually initialized the parameters of control policy with pre-computed gains from state-feedback controllers, e.g. Linear-Quadratic Regulators, which requires user time and control expertise to properly perform the initialization process, ANLC instead removes the need of control initialization by its proposed compositional control architecture containing both linear and nonlinear control laws so that the proposed method allows the synthesis of nonlinear (as well as linear) control laws with the sole requirement being the knowledge of the system dynamics. For empirical experiments, we tested the ANLC algorithm for all system dynamics in Appendices F and G. We tested on the Van Der Pol Oscillator and 3D Trig dynamics to get the best hyperparameter setting. In bold, we show the chosen parameters, selected to have the best success discovery rate on Van Der Pol Oscillator and 3D Trig Dynamics.
@@
FOSSIL 2.0 (Edwards et al., 2024) is a software tool for robust formal synthesis of certificates (e.g., Lyapunov and barrier functions) for dynamical systems modelled as ordinary differential and difference equations. FOSSIL 2.0 implements a counterexample-guided inductive synthesis (CEGIS) for the construction of certificates alongside a feedback control law. In the loop of CEGIS, the learner, based upon neural network templates, acts as a candidate to satisfy the conditions over a finite set $\mathcal{D}$ of samples, while the verifier (formal verification tools) works in a symbolic environment that either confirms or falsifies whether the candidate from learner satisfies the conditions over the whole dense domain $\mathcal{X}$. If the verifier falsifies the candidate, one or more counterexamples identified by the verifier are added to the sample set, and the network is retrained. This loop repeats until the verification proves that no counterexamples exist or until a timeout is reached. Similar to the ANLC, in the empirical experiment, we set the hyperparameters based on the Van Der Pol Oscillator and 3-D Trig dynamics and tested for all other dynamics in Appendices F and G.
@@
E.3 Global Lyapunov Function Discovery by Pre-trained Transformer
Alfarano et al. (2024) pre-trained a transformer on backward-generated and forward-generated global Lyapunov function datasets. The backward-generated datasets involve sampling arbitrary positive definite functions and deriving corresponding stable dynamics through some specific symbolic designs, while the forward-generated polynomial datasets contain randomly generated dynamics with corresponding Lyapunov functions identified by SOS methods if the system is inherently globally stable. Candidate Lyapunov expressions are sampled using beam search in a token-by-token manner. However, their method cannot adaptively refine the candidate Lyapunov functions if the beam search fails on specific dynamics, and it requires a dataset that is expensive to generate (e.g., thousands of CPU hours for a 5-D dynamics dataset) to achieve adequate generalization during inference. Furthermore, its emphasis on global stability limits its applicability to real-world, nonpolynomial control systems, which typically only admit local stability. Due to the lack of resources of multiple industrial-level GPUs, we contacted the authors of Alfarano et al. (2024) to conduct the evaluation of their pre-trained model on our test systems, which is shown in Section 5.
E.4 Sum-of-Squares (SOS) Methods
SOS methods formulate Lyapunov functions discovery of given dynamics as a semi-definite programming task, where the coefficients of a pre-defined SOS candidate expressions are optimized to satisfy the Lyapunov conditions (hard constraints in the optimization problem) using convex optimization tools. SOS methods are generally applied to polynomial systems for stability analysis. With proper recasting techniques, SOS methods can also be applied to non-polynomial systems.
Definition E.1 (Sum of Squares, Papachristodoulou & Prajna (2005a)).
For $x\in\mathbb{R}^{n}$, a multivariate polynomial $p(x)$ is a sum of squares (SOS) if there exist some polynomials $f_{i}(x),i=1,\cdots,M$ such that
$p(x)=\sum_{i=1}^{M}f_{i}(x)^{2}.$
[…]
  1. 1.
    $V(x)-\sum_{i=1}^{n}\sum_{j=1}^{d}\epsilon_{ij}x_{i}^{2j}$ is a SOS, where $\sum_{j=1}^{d}\epsilon_{ij}>\gamma,\forall i=1,...,n$ with $\gamma>0$, and $\epsilon_{ij}\geq 0$ $\forall$ $i$ and $j$,
  2. 2.
    $-\frac{\partial V}{\partial x}f(x)$ is a SOS.
For local stability analysis, consider a ball of radius $r$ centered at origin $\mathcal{B}_{r}(0)$, which can be represented by the semialgebraic set $S=\{x:g(x,r)\geq 0$, where $g(x,r)=r-\sum_{i=1}^{n}x_{i}^{2}\}$. We require that the stability condition holds in $S$. Retaining the same optimization objective and constraints on $V(x)$ as before, a modified constraint on Lie derivative is imposed: $-\frac{\partial V}{\partial x}f(x)-s(x)g(x,r)$ is a SOS for some SOS $s(x)$. If such an $s(x)$ exists, we can establish local stability.
In Section 5, we develop our code based on the findlyap function from SOSTOOLS (MATLAB) and issue-16 of SOSTOOLS’ official GitHub repo to examine the SOS method on polynomial systems in Appendix F. Table 5 summarizes the experiment results of the SOS approach on our polynomial test dynamics.
[…]
Systems App. F.1 App. F.2 App. F.3 - I App. F.3 - II App. F.4 App. F.5 App. F.6
Degree 2d 2 2 2 4 2 2 2
Region $\mathcal{B}_{1}(0)$ Global Global Global $\mathcal{B}_{1}(0)$ $\mathcal{B}_{1}(0)$ $\mathcal{B}_{1}(0)$
Runtime 0.697s 0.832s 0.497s 2.509s - - -
Non-Polynomial System. To apply SOS method on non-polynomial system $\dot{z}=f(z),z\in\mathcal{D}_{1}$, define $x_{1}=z$ as the original states and $x_{2}$ as the newly introduced variables to recast non-polynomial terms in $f(z)$. Let $x=(x_{1},x_{2})$. The system dynamics can then be written in rational polynomial forms:
$\begin{cases}\dot{x}_{1}&=f_{1}(x),\\ \dot{x}_{2}&=f_{2}(x),\end{cases}$
with constraints $x_{2}=F(x_{1}),G_{1}(x)=0,G_{2}(x)\geq 0,$ where $F,G_{1},G_{2}$ are vectors of functions, to restrict the states of recast dynamics $x=(x_{1},x_{2})$ equal to the manifold of the original state $z$.
Define $g(x)$ as the collective denominator of $f_{1},f_{2}$, and local region of interest as a semialgebraic set:
$\{(x)\in\mathbb{R}^{n+m}|G_{\mathcal{D}}(x)\geq 0\},$
where $G_{\mathcal{D}}(x)$ is a vector of polynomials designed to match the original state space. Let $x_{2,0}=F(0)$. Suppose there exist polynomial functions $V(x)$, $\lambda_{1}(x),\lambda_{2}(x)$, and SOS polynomials $\sigma_{i}(x),i=1,2,3,4$, of appropriate dimensions such that
$V(0,x_{2,0})=0,$ (10)
$V(x)-\lambda_{1}^{T}(x)G_{1}(x)-\sigma_{1}^{T}(x)G_{2}(x)-\sigma_{3}^{T}(x)G_{D}(x)-\phi(x)\in\text{SOS},$ (11)
$-g(x)(\frac{\partial V}{\partial x_{1}}(x)f_{1}(x)+\frac{\partial V}{\partial x_{2}}(x)f_{2}(x))-\lambda_{2}^{T}(x)G_{1}(x)-\sigma_{2}^{T}(x)G_{2}(x)-\sigma_{4}^{T}(x)G_{D}(x)\in\text{SOS},$ (12)
where $\phi(x)$ is some scalar polynomials with $\phi(x_{1},F(x_{1}))>0,\forall x_{1}\in\mathcal{D}_{1}\backslash\{0\}$. If constraints (10), (11), and (12) hold, then $z=0$ is stable (not necessarily asymptotically stable) (Papachristodoulou & Prajna, 2005b). In Section 5, we develop our code from SOSTOOLS to examine the SOS methods on two non-polynomial systems: simple pendulum and 3-D trig dynamics (Appendices G.1 & G.2).
@@ F.3 Three-variable-polynomial-systems with higher degree
Table 6: Three-Dimensional Polynomial Example with Higher Degree.
@@
Our proposed method is trained over the state space $\mathcal{D}=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{6}\bm{\mid}|x_{i}|\leq 1,\forall i\in\{1,2,..,6\}\}$, and is able to find a valid Lyapunov function $V(x)=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2}+x_{6}^{2}$. This dynamics is not globally asymptotically stable since if $x_{1},x_{2},\text{ or }x_{5}$ has a significantly large value, the perturbations introduced by the small gains will shift the object by a significant amount away from the equilibrium point. By empirical checking, our found Lyapunov function certifies the asymptotical stability of this system over the region $\mathcal{D}^{\prime}=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{6}\bm{\mid}\sum_{i=1}^{6}x_{i}^{2}\leq 500\}$.
@@ G.3 N-bus Lossless Power System
We test our proposed framework on the N-bus power lossless system (Cui et al., 2023a; Feng et al., 2024a) to examine its ability to handle complex high-dimensional dynamics. Consider $\theta_{i}$, $\omega_{i}$ as the phase angle and the frequency of bus $i$, respectively, the dynamics for each bus are formulated as follows,
@@
where $\delta_{i}$ can be understood as the center-of-inertia coordinates of each bus. In our experiment, we test the proposed framework on the 3-bus power system. For simplicity, we set $p_{i}=0$, $m_{i}=2$, $d_{i}=1$, $u_{i}(\omega_{i})=\omega_{i}$, and $B_{ij}=1\>\forall\>i\neq j,B_{ii}=0$. In this case, the equilibrium point for our system is at the origin, i.e. $\delta_{i}^{*}=\omega_{i}^{*}=0,\>i=1,2,3$. The state space for our experiment is defined as: $\mathcal{D}=\{(\delta_{1},\delta_{2},\delta_{3},\omega_{1},\omega_{2},\omega_{3})\in\mathbb{R}^{6}\bm{\mid}|\delta_{i}|\leq 0.75\text{ and }|\omega_{i}|\leq 1.2\text{ for }i=1,2,3\}$. Through our method, we retrieved a valid Lyapunov function $V(\delta_{1},\delta_{2},\delta_{3},\omega_{1},\omega_{2},\omega_{3})=(\sum_{i=1}^{3}\omega_{i}^{2})-0.5\left(\sum\limits_{i=1}^{3}\sum\limits_{\begin{subarray}{c}j=1,i\neq j\end{subarray}}^{3}\cos(\delta_{i}-\delta_{j})-1\right)$, which is consistent to the known Lyapunov function presented in Cui et al. (2023a). The Lie derivative of the identified Lyapunov function can be simplified as $L_{f}V=-2(\omega_{1}^{2}+\omega_{2}^{2}+\omega_{3}^{2})$. The analytical structure of this found Lyapunov function and invariance principle allows us to easily identify it as a valid Lyapunov function by hand.
@@ G.4 Indoor Micro Quadrotor
For the angular rotations subsystems of the quadrator from Bouabdallah et al. (2004), it has 6 states to describe the angular motion of the quadrotor. The states $x_{1},x_{3},\text{ and }x_{5}$ describe the roll, pitch, and yaw of the quadrator, and states $x_{2},x_{4},\text{ and }x_{6}$ represent their time derivatives. With perturbation terms $\Omega$ and control inputs $U_{1},U_{2},\text{ and }U_{3}$, the subsystem can be formulated as follows,
@@
and restricting $I_{x}=I_{y}$, the angular rotations subsystems is stabilized to the chosen equilibrium point $X_{d}=\{x_{1}^{d},0,x_{3}^{d},0,x_{5}^{d},0\}$. In empirical experiments, we set $X_{d}=\{0,0,0,0,0,0\}$, state space $\mathcal{D}=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{6}\bm{\mid}|x_{i}|\leq 3,\forall i\in\{1,2,..,6\}\}$, $I_{x}=I_{y}=2$, $I_{z}=5$, $k_{1}=5$, $k_{2}=20$, $k_{3}=4$, $l=1$, $J_{R}=1$, and $\Omega=\sin(x_{2})\cos(x_{4})$, our framework successfully found Lyapunov function $V(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=\sum_{i=1}^{6}x_{i}^{2}$. By examining the Lie derivative $L_{f}V=-2.5x_{2}^{2}-10x_{4}^{2}-0.8x_{6}^{2}$ and using the Invariance principle, we conclude that this subsystem is globally asymptotically stable. Under other parameter settings, our framework can also retrieve valid analytical Lyapunov functions for this dynamics.